OpenStudy (anonymous):

Hi everyone! For this parametric, find dy/dx x=sinh(t) y=cosh(t)...My final answer was tanh(t)...can anyone confirm that for me? Thanks! :o)

jimthompson5910 (jim_thompson5910):

x = sinh(t) dx/dt = cosh(t) y = cosh(t) dy/dt = sinh(t) ------------------------ dy/dx = (dy*dt)/(dt*dx) dy/dx = (dy/dt)*(dt/dx) dy/dx = (dy/dt)/(dx/dt) dy/dx = sinh(t)/cosh(t) dy/dx = tanh(t) So it looks good to me

OpenStudy (anonymous):

thanks! @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yw